# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

### v2

@@ V-C3 Evaluation Methods
In the ablation study on model training methods, we aimed to explore how different training approaches impact model performance. We compared three training methods: (1) the method used in our benchmark, which selects the epoch with the best performance on the validation set (as “Ours” in Tables VII and VIII); (2) selecting the model’s result from the last epoch (as “R1” in Tables VII and VIII), where, for a fair comparison, the model’s training set includes the validation set; and (3) selecting the epoch when the loss value on the validation set decreases by less than 1% within the first 10 epochs (as “R2” in Tables VII and VIII).
@@ VI Future Direction
Experimental Settings. LibEER includes several key experimental setups to facilitate research, including four widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods. However, LibEER does not cover all experimental configurations used in the EER field. Some studies use other datasets like DREAMER[47], Amigos[48], and MPED[49]. Others develop unique preprocessing methods, including customized filtering, artifact removal[8, 52, 53], denoising[50, 51], normalization techniques[51, 8], and extracting other types of features [12, 45, 15]. Additionally, some tasks, such as cross-session[54, 55, 56], subject-independent[55, 38], and cross-dataset tasks[46, 51], are designed to assess model performance under different conditions, such as data from different sessions or populations. Furthermore, alternative evaluation strategies, including early stopping[57, 58, 59], are employed in some studies for more robust model training and assessment. Future versions of LibEER will aim to incorporate more diverse experimental setups from the EER field.
@@ VII Conclusion
In this work, we propose LibEER, a comprehensive benchmark and algorithm library specifically designed for EEG-based emotion recognition. By standardizing the experimental settings, including dataset usage, data preprocessing methods, experimental task selection, data splitting strategies, and evaluation methods, LibEER addresses the prevalent issues of inconsistency and reproducibility in the EER field. Through its unified PyTorch codebase and the comprehensive implementation of over ten representative models, our benchmark ensures fair comparisons and accurate evaluations of models in both subject-dependent and cross-subject scenarios. Additionally, through fair comparison experiments and ablation studies, we present findings and discussions that aim to inspire researchers in the field. We believe that LibEER will not only facilitate the development and benchmarking of new EER models but also promote standardization and transparency, thereby advancing research and applications in this domain. In the future, we plan to further enhance LibEER by incorporating additional experimental settings and more advanced models.

### v3

@@ V-C3 Evaluation Methods
In the comparative experiments on model training methods, we explored how different training approaches impact model performance. We compared three training methods: (1) the method used in our benchmark, which selects the epoch with the best performance on the validation set (as “Ours” in Tables VII and VIII); (2) selecting the model’s result from the last epoch (as “R1” in Tables VII and VIII), where, for a fair comparison, the model’s training set includes the validation set; and (3) selecting the epoch when the loss on the validation set decreases by less than 1% within the first 10 epochs (as “R2” in Tables VII and VIII).
@@
V-C4 Random Seeds
The random seed determines the reproducibility of experimental results, yet it is often unspecified in existing studies. To investigate the impact of random seed settings on model performance, we conducted comparisons using seeds set to 2023, 2024, 2025, 2026, and 2027. The results are shown in Figure 9. Two key conclusions can be drawn from the results. First, the random seed has a significant impact on experimental outcomes, with different models exhibiting considerable performance variation under different seeds. This is because changing the random seed affects both the random data splitting process and the stochastic behaviors during model training, such as parameter initialization, mini-batch ordering, and dropout, all of which can lead to variations in the final results. Among all the methods, CDCN is most affected by the random seed, with performance differences reaching approximately 20%. In contrast, HSLT and ACRNN are relatively less affected, though their variations still reach around 10%. Second, in the majority of cases, these models exhibit similar variation patterns when the random seed changes. This suggests that using a unified random seed across all models can help ensure the fairness of comparisons to some extent. This finding underscores the importance of adopting the same random seed for all methods in the EER field and explicitly reporting the seed value in publications to ensure fairness and reproducibility of results.
@@ VI Future Direction
Experimental Settings. LibEER includes several key experimental setups to facilitate research, including six widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods. However, LibEER does not cover all experimental configurations used in the EER field. Some studies use other datasets like DREAMER[54] and Amigos[55]. Others develop unique preprocessing methods, including customized filtering, artifact removal[11, 58, 59], denoising[56, 57], normalization techniques[57, 11], and extracting other types of features [15, 52, 18]. Additionally, some tasks, such as cross-session[60, 61, 62], subject-independent[61, 45], and cross-dataset tasks[53, 57], are designed to assess model performance under different conditions, such as data from different sessions or populations. Furthermore, alternative evaluation strategies, including early stopping[63, 64, 65], are employed in some studies for more robust model training and assessment. Future versions of LibEER will aim to incorporate more diverse experimental setups from the EER field.
@@ VII Conclusion
In this work, we propose LibEER, a comprehensive benchmark and algorithm library specifically designed for EEG-based emotion recognition. By standardizing the experimental settings, including dataset usage, data preprocessing methods, experimental task selection, data splitting strategies, and evaluation methods, LibEER addresses the prevalent issues of inconsistency and reproducibility in the EER field. Through its unified PyTorch codebase and the comprehensive implementation of seventeen representative models, our benchmark ensures fair comparisons and accurate evaluations of models in both subject-dependent and cross-subject scenarios. Additionally, through fair comparison experiments and investigation of key experimental settings, we present findings and discussions that aim to inspire researchers in the field. We believe that LibEER will not only facilitate the development and benchmarking of new EER models but also promote standardization and transparency, thereby advancing research and applications in this domain. In the future, we plan to further enhance LibEER by incorporating additional experimental settings and more advanced models.